OpenStudy (anonymous):

Find the area to the left of 39.213 under the X^2 distribution.

OpenStudy (anonymous):

What are the boundaries?

OpenStudy (anonymous):

Are you integrating?

OpenStudy (anonymous):

I honestly am not sure how to answer your questions. This whole chapter is really confusing me :(

OpenStudy (anonymous):

I can help you solve one similar to that.

OpenStudy (anonymous):

That works for me. I just need to know how to do it

OpenStudy (anonymous):

An insurance agent says the standard of the total hospital charge for patients involved in a crash in which the vehicle struck a construction barricade is less than $3500. A random sample of 28 total hospital charges for patients involved in this type of crash has a standard deviation of $4100. at a = 0.10, can you support the agent's claim? Solution: The test hypothesis is Ho:σ=3500 Ha:σ<3500 The test statistic is chisquare= (n-1)*s^2/σ^2 =27*4100^2/3500^2 =37.05061 Given a=0.1, the critical value is chisquare with 0.99, df=n-1=26 is 45.64 (check chisquare table) Since 37.05061 is less than 45.64, we do not reject Ho. So we can not support the agent's claim

OpenStudy (anonymous):

I've been looking at this problem for a little bit now and I still don't understand at all how to get the P value. Then based off of the value you got off the chisquare table I was still wrong. It was .9394.. :/

OpenStudy (hba):

@KiwiBaby Are you done ?

OpenStudy (anonymous):

Nope, not at all, I just have to start all over on a question just like it and I will be stuck on the same place next time :(

OpenStudy (hba):

@KiwiBaby Do you mind going through the basics again ?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

@hba Are you still there?

OpenStudy (hba):

I am really sorry i am working on a tutorial for you guys,maybe @waterineyes can help here :)

OpenStudy (anonymous):

That's ok. I somehow managed to figure it out! I'm thrilled.. Thank you guys.